OpenStudy (anonymous):

Solve for x: log(x^6)=log(x)^2 Note that there are 2 solutions. A & B where A

OpenStudy (anonymous):

e^6

hartnn (hartnn):

use this: \(\huge log b^n=nlogb\)

OpenStudy (anonymous):

\[ \log{x^6}=(\log{x})^2\\ 6\log x=\log{x}\log{x}\\ 6=\log{x}\\ x=10^6 \]This presumes that this is the common logarithm.

OpenStudy (anonymous):

i assumed that it is in e base